OpenStudy (anonymous):

How would I graph: -3(x-2)^3 + 24

OpenStudy (anonymous):

See the attachment.

OpenStudy (anonymous):

But how do I do that? I can't use a calculator on the test @robtobey

OpenStudy (anonymous):

here u gooo

OpenStudy (anonymous):

I don't really get it

OpenStudy (anonymous):

u have to change the windows in calculator

OpenStudy (anonymous):

I cant use the calculator on the test tomorrow........

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

y = -3(x-2)^3 + 24 As you can see this is a cubic curve. You will have to know what the general y = x^3 curve and y = -x^3 curve looks like. Here the curve will look more like the y = -x^3 curve just shifted to the right by 2 units, shifted up by 24 units and vertically stretched by a factor of 3. First find the y-intercept by setting x = 0 Then find the x-intercept by setting y = 0.

OpenStudy (ranga):



OpenStudy (ranga):